Find the first 25 terms of the arithmetic series:
8+15+22+29+...

Answer by partha_ban(41)

We know that an = a1 + (n – 1)d
where an = n-th term = 25th term here
a1 = 1st term = 8
d = constant difference = 15-8 = 7
n = 25
Therefore, a25 = 8 + (25-1)*7
= 8 + 24*7
= 8 + 168
= 176